# ELMS: Elasticized Large Language Models On Mobile Devices

### v1

@@ 5.2. End-to-end performance
We first evaluate end-to-end performance on traces in §$\S$5.1.
@@
ELMS significantly outperforms the baselines by 6.00%–16.83% (11.04% on average) in absolute accuracy. Compared to PFS, ELMS further involves prompt elastification and does not introduce costly switching overhead. Besides, PFS does not fully make use of the room below a given SLO. Another potential reason is that ELMS derives sub-models from a bulky one, which may feature stronger emergent ability than the SLMs pretrained from scratch. Compared to LPruner, ELMS involves prompt elastification, and minimizes switching overhead. Compared to LE, ELMS employs fine-grained permutation consistent units instead of the coarse-grained entire layers for elastification. Although both ELMS and LG2+CS orchestrate prompt- and model- elastification, there are many essential issues that degrade the performance of LG2+CS. Firstly, the heavy SLM of LLMLingua2 must compress the prompt very aggresively to meet a tight TTFT SLO since contextual sparsity shows limited acceleration on prefill stage due to the low locality. Secondly, the sparsity ratio is also compromised due to non-relu (Agarap, 2019; Touvron et al., 2023) activation functions, non-elastictizable attention (Liu et al., 2023) and degraded performance of sparse kernels.
@@
Figure 16. Online overhead. “LP” means LPruner.
Memory consumption. We discuss peak memory consumption in Figure 16. Without loss of representativeness. we report LLaMA-7B in the trace with $\alpha=0$ on Redmi K60 Champion. ELMS consumes on-par memory compared to the baselines (15–17GB). Notably, directly deploying all the dedicated size LLMs in memory is impractical. As marked as PFS(Ideal) in Figure 16, it consumes 29.3GB memory in total, which is OOM on all the COTS devices in Table 2.
Switching overhead. With the same setting, we report the breakdown latency of switching between different model elastification levels (i.e., sub-models) in Figure 16. Switching time is actually part of TTFT, and a too long switching time will preempt the room of model/prompt capacity, leading to a lower accuracy. PFS (Swap) and LPruner incur unacceptable time overhead that up to 8.3/6.2 seconds per request. This is mainly attributed to the costly swapping and in-memory data movement. ELMS only takes 0.31 second to switch to a new sub-model. Such a number is lower than 1% of LLMaaS’s average TTFT, thus being completely acceptable.
Remarks ELMS is the most high-quality and feasible solution for end-to-end on-device elastic LLM service.
@@ 5.3. Performance on standalone datasets
With a specific SLO, we further report the performance on an entire standalone dataset to show ELMS’s superiority. The switching overhead is dismissed as there is no upgrade/downgrade. The results are obtained on cloud server with SLO statistics on Mi 14 smartphone, and are shown in Figure 15. We have the following observations. ELMS significantly outperforms all its baselines by up to 34% and on-average 22% on accuracy. Specifically, on all the SLOs, ELMS always provides a much higher accuracy than the baselines (LE/LG2+CS/LPruner) that derive sub-models from the original one. Compared to PFS, ELMS provides a higher accuracy on 88.4% of all SLOs, and the maximal accuracy loss is within 9.6%. The reasons are as discussed before. Please also note that PFS is costly in terms of switching and memory at runtime.
@@ 5.4. Offline stage overhead
Figure 17. Offline overhead of ELMS.
We further analysis the offline overhead in Figure 17. We measure the elastification of LLaMA-7B for MI14 on our A40 cloud server. Compared to PFS that trains a dedicated LLM for each SLO, ELMS derives elasticized LLMs from the original LLaMA-7B. Thereby, as shown in Figure 17, the entire offline stage of ELMS only consumes 68.3 GPU hours (translates to about \$100 GPU renting price (hfg, 2024)), making it affordable for most LLMaaS developers. Compared to other baselines that also derive elasticized LLMs from the original one, ELMS takes 21.7–68.2 more GPU hours, which is acceptable since the offline stage is performed once for all. Note that in the above sections, we have demonstrated that ELMS delivers an elastic LLM service with much higher quality than these baselines.
In Figure 17, we provide a detailed breakdown of ELMS’s offline stage. The model recovery (§$\S$3.2) and self-induced labelling (§$\S$3.3) dominate the offline stage, taking 40.4/26.3 hours. The reason the latter requires tremendous time is due to the lower GPU utilization caused by interactions with the score-head and sub-models.
@@ 5.5. Sensitivity analysis
Data sensitivity. As shown in Figure 18, the data scale of elastification exhibits a marginal effect. Regarding to model recovery, we further collect training data from LaMini (Wu et al., 2023) dataset that akin to Alpaca-cleaned. The final accuracy only increases 2.9%/3.7% when the recovery corpus is 10$\times$/20$\times$ larger. Regarding to unit importance profiling and decision training data, we expand them from Bookcorpus and MMLU (Hendrycks et al., 2021), respectively. We also observe the similar marginal effect. Since Meetingbank is currently the largest corpus for token importance scoring to the best of our knowledge, we leave expanding it as a further work. In a nutshell, ELMS’s data scale achieves a strong and sweet spot for high-quality elastification.
Figure 18. Data sensitivity. On trace $\alpha=0$, LLaMA-7B.
Configuration sensitivity. We discuss the sensitivity of ELMS’s configurations in Figure 19. In Figure 19, we use the trace of $\alpha=0$ on LLaMA-7B. We have the following conclusions. (1) 40M parameters are already a sweet config. of TLM scale. A larger TLM’s gain on scoring and decision-making accuracy quickly gets eliminated by the overhead. (2) 12 shared bottom layers are reasonable for TLM, since more will lead to a smaller capacity of the heads, and less will incur a higher inference/training overhead. (3) A step size of 10% for model and prompt elastification is fine-grained enough for serving the diversified SLOs. Shrinking it to 5% makes almost no difference. We also provide a more comprehensive SLO configuration in Figure 19 and report the delta accuracy of ELMS compared to its strongest baseline PFS (without considering switching overhead) on ARC_E LLaMA-7B. ELMS is close to or much better than PFS on all SLOs, especially those that have a much more tighter TTFT and are more common on mobile LLM tasks. In end-to-end traces, ELMS can even achieve a better performance with its overhead-free switching.
(a) I. TLM size. II. # of shared layers. III. Elasticity stepsize.
(b) More SLOs.
Figure 19. Configuration sensitivity.
@@
Quantization. Quantization is a widely-used technique for deploying LLMs on mobile devices. Here we employ a linear and weights only method on the $\alpha=0$ trace in Figure 20. ELMS can deliver LLM service with almost lossless accuracy under 8bits integers, and acceptable (3% lower absolute accuracy) under 4bits.
Concise prompts. We further emulate the scenario where the prompts have been consciously streamlined by the LLMaaS callers. We use LLMLingua2 to filter out about 15% verbose tokens in the prompt. The results (on trace with $\alpha$=0) in Figure 20 show that ELMS can still significantly outperform its baselines.
5.6. Generalization to other SLOs
@@
Although in this paper we identify LLM inference latency as the SLO of LLMaaS requests, the SLO can also been easily generalized to other metrics with our proposed elastification methods. Here we define another two SLOs: end-to-end request inference time $SLO_{time}$ (i.e., prefill + decode stage) and inference energy consumption $SLO_{energy}$, each of which is with a step size of 10% full LLMaaS. The results are obtained on Mi 14 smartphone. As shown in Figure 21, ELMS consistently outperforms its baselines on these new SLOs. the rationale is that these SLOs are all a specific kind of resource requirement and can be break down into prompt- and model- level elastifications. We believe that ELMS can serve various metrics of SLOs for diverse apps’ requirements.
5.7. On mobile SoC accelerators
@@
Figure 22. “What if” experiment on mobile accelerators that cannot seamlessly run ELMS and its baselines currently. LG2+CS is omitted due to its sparse compute.
Mobile devices (smartphones) are typically equipped with hardware accelerators (e.g., Adreno GPUs or Qualcomm NPUs) in their SoCs. These SIMD/SIMT DSAs achieve higher throughput and lower latency on traditional on-device DNNs, especially CNNs. Yet, to the best of our knowledge currently (Aug., 2024) none of them can seamlessly run prefill and decode of ELMS and all its baselines due to legacy SDK/hardware issues like dynamic shape, graph building or non-sparse kernels. Thus, we conduct a “what if” experiment by mapping the theoretical compute/memory load to the profiled latency on these processors. We use MNN (Jiang et al., 2020) and mllm-NPU (Xu et al., 2024b) for GPU/NPU profiling on Mi 14, respectively. The results are shown in Figure 22. ELMS can still significantly outperform the baselines under this setting. We believe that ELMS will become more friendly and practical to practitioners with the maturity of accelerators.
@@ 6. Related Work
Elastic neuron networks. Elastic neuron networks can change their capacity at runtime to dynamically adapt to various resource/accuracy constraints. Early exit networks (Teerapittayanon et al., 2017; Tambe et al., 2021) only perform inference on bottom layers since they are empirically more critical. Yet, early exit is not suitable for elastic LLM service. On one hand, due to LLMs’ autoregressive inference nature, a skipped layer’s KV cache may be accessed later. On the other hand, using layers as the granularity for trade-offs is not fine-grained enough. Parameter sharing networks (Fang et al., 2018; Han et al., 2021; Wen et al., 2023) generate memory-efficient sub-models that shares parameters with each other. For instance, NestDNN (Fang et al., 2018) and LegoDNN (Han et al., 2021) create sub-models of CNNs via offline re-pretraining, which is costly for foundation models. Adaptivenet (Wen et al., 2023) employs a CNN-oriented supernet, which provides diverse accuracy-latency trade-offs yet needs extensive pre-training and memory resources. Activation sparsity (Liu et al., 2023; Song et al., 2023; Cao et al., 2019; Kong et al., 2023) elasticitizes DNNs via sparsifying weights according to the NN inputs. An accuracy-latency trade-off can be achieved by setting the proportion of activated weights. Yet, they either focus on CNNs with relu (Agarap, 2019) structure (e.g., convrelu++ (Kong et al., 2023), seernet (Cao et al., 2019)), or only accelerates the decoding stage of relu-based MLPs of LLMs. The LLM prefill stage cannot be elastictized due to the low locality. And the trade-off on decode stage is limited by both the sparse kernels and the attention module. Besides, the non-relu LLMs (e.g., LLaMA (Touvron et al., 2023), Vicuna (Zheng et al., 2023a)) exhibit a poor performance without resource-extensive pre-training. In a nutshell, ELMS outperforms the above work by employing an orchestrated prompt-model elastification, providing a high-quality accuracy-latency trade-off in both TTFT and TPOT for LLMaaS apps.
@@
Efficient on-device LLM inference. Tremendous work (team, 2023; Jiang et al., 2020; mll, 2024; Xue et al., 2024; Xu et al., 2023; Yi et al., 2023; Xu et al., 2024a) shed light on resource-efficiently deploying LLMs on mobile devices. For instance, MLC-LLM (team, 2023) is an NN-compiler with operator- and kernel- level optimizations for mobile devices. MNN (Jiang et al., 2020) and mllm (mll, 2024) are on-device inference libraries for LLMs. PowerinferV2 (Xue et al., 2024) addresses the memory issue of mobile devices by introducing swapping and activation sparsity to LLM inference. Targeting at elastic LLM service, ELMS is orthogonal to these work, and can benefit from them to realize the vision of on-device LLMaaS.
@@
Foundation models as a service. As a general task solver, or so-called “AGI”, a single foundation model is deployed as a service to process heterogeneous tasks (Yin et al., 2024; Yuan et al., 2024; Ge et al., 2023). For instance, AIOS (Ge et al., 2023) is a system where a single LLM runs as the “brain” to serve all apps on the device. ELMS makes it more practical on resource-constrained mobile devices.
@@
Model collaboration. ELMS employs a dual-head tiny language model to elasticize the prompt of the LLM. Using a small model to collaborate with the big model is common in ML systems. For instance, speculative decoding (Miao et al., 2024; Xu et al., 2023) accelerates the decode stage with a draft SLM. However, the prefill stage cannot be accelerated since it is typically compute-bound. LLMlingua (Jiang et al., 2023) use an SLM to refine the LLM prompt, which has several untackled issues for elastic LLM service as discussed in this paper. ELMS’s TLM is effective an efficient for elastic on-device LLMaaS.
@@ 7. Conclusion
This work has proposed ELMS, the first-of-its-kind elastic on-device LLM service that serves apps with diverse SLOs. ELMS incorporates two novel designs, i.e., one-shot reordering of permutation consistent units and dual-head tiny language model to fully unleash the potential of model- and prompt- elastification. ELMS significantly outperforms competitive baselines by up to 16.83% and 11.04% on average in accuracy.

### v2

@@ 5.2. End-to-end performance
We first evaluate end-to-end performance on traces in $\S$5.1.
@@
ElastiLM significantly outperforms the baselines by 6.60%–14.83% (10.45% on average) in absolute accuracy. Compared to PFS, ElastiLM further involves prompt elastification and does not introduce costly switching overhead. Besides, PFS does not fully make use of the room below a given SLO. Another potential reason is that ElastiLM derives sub-models from a bulky one, which may feature stronger emergent ability than the SLMs pretrained from scratch. Compared to LPruner, ElastiLM involves prompt elastification, and minimizes switching overhead. Uniqueness to LG2+CS. Firstly, the heavy SLM of LLMLingua2 must compress the prompt very aggresively to meet a tight TTFT SLO since contextual sparsity shows limited acceleration on prefill stage due to the low locality. Secondly, the sparsity ratio is also compromised due to non-relu (Agarap, 2019; Touvron et al., 2023) activation functions, non-elastictizable attention (Liu et al., 2023) and degraded performance of sparse kernels. In contrast, ElastiLM’s model elastification identifies and leverages the permutation consistency in Transformer models and works for both prefill and decoding, and its prompt elastification considers and tackles the unique challenge of prompt-model orchestration. Compared to LE/LaCo/ShortGPT/AttnDrop, ElastiLM shows steady performance gain. This is due to that these methods only elasticizes the model in layer level. In contrast, ElastiLM’s unit-level fine-grained pruning traverses a much larger space. Also, LaCo will generate sub-models that cannot fully share the weights between each other, resulting in extra switching overhead.
@@
Figure 16. Online overhead.
Memory consumption. We discuss peak memory consumption in Figure 16. Without loss of representativeness. we report LLaMA-7B in the trace with $\alpha=0$ on Redmi K60 Champion. ElastiLM consumes on-par memory compared to the baselines (15–17GB). Notably, directly deploying all the dedicated size LLMs in memory is impractical. As marked as PFS(Ideal)/LaCo(Ideal) in Figure 16, it consumes 29.3GB/76.3GB memory in total, which is OOM on all the COTS devices in Table 2.
Switching overhead. With the same setting, we report the breakdown latency of switching between different model elastification levels (i.e., sub-models) in Figure 16. Switching time is actually part of TTFT, and a too long switching time will preempt the room of model/prompt capacity, leading to a lower accuracy. PFS (Swap) and LPruner incur unacceptable time overhead that up to 8.3/6.2 seconds per request. This is mainly attributed to the costly swapping and in-memory data movement. ElastiLM only takes 0.31 second to switch to a new submodel. Such a number is lower than 1% of average TTFT of LLM service, thus being completely acceptable. The end-to-end experiments with switching overhead considered in Figure 14 also show that the switching time of ElastiLM is an acceptable tradeoff for LLM service.
Remarks ElastiLM is the most high-quality and feasible solution for end-to-end on-device elastic LLM service.
@@ 5.3. Performance on standalone datasets
With a specific SLO, we further report the performance on an entire standalone dataset to show ElastiLM’s superiority. The switching overhead is dismissed as there is no upgrade/downgrade. The results are obtained on cloud server with SLO statistics on Mi 14 smartphone, and are shown in Figure 15. We have the following observations. ElastiLM significantly outperforms all its baselines by up to 40% on accuracy. Specifically, on all the SLOs, ElastiLM always provides a much higher accuracy than the baselines (LE/LG2+CS/LPruner) that derive sub-models from the original one. Compared to PFS, ElastiLM provides a higher accuracy on 77.8% of all SLOs. The reasons are as discussed before. Please also note that PFS and LaCo are costly in terms of switching and memory at runtime. Performance gains are minimal in some high-SLO scenarios, this is due to a slight pruning (i.e., high SLO) will not significantly differentiate ElastiLM from the pruning baselines. However, only ElastiLM can work for all the varying SLOs for apps, since there will not always be high-SLO scenarios, and switching overhead (Figure 16) will also be considered when serving multiple apps.
@@ 5.4. Offline stage overhead
Figure 17. Offline overhead of ElastiLM.
(a) I. Model Recovery. II. Importance profiling. III. Decision-head training.
(b) I. TLM size. II. # of shared layers. III. Elasticity stepsize.
(c) Ablation study.
(d) Varying real-world SLOs.
Figure 18. Sensitivity analysis and ablation study. (a)/(b)/(c): LLaMA-7B, on trace with $\alpha$=0; (d): LLaMA-7B, ARC_E.
We further analysis the offline overhead in Figure 17. We measure the elastification of LLaMA-7B for MI14 on our A40 cloud server. Compared to PFS that trains a dedicated LLM for each SLO, ElastiLM derives elasticized LLMs from the original LLaMA-7B. Thereby, as shown in Figure 17, the entire offline stage of ElastiLM only consumes 68.3 GPU hours (translates to about \$100 GPU renting price (hfg, 2024)), making it affordable for most developers. Compared to other baselines that also derive elasticized LLMs from the original one, ElastiLM takes 21.7–68.2 more GPU hours, which is acceptable since the offline stage is performed once for all. Note that in the above sections, we have demonstrated that ElastiLM delivers an elastic LLM service with much higher quality than these baselines.
In Figure 17, we provide a detailed breakdown of ElastiLM’s offline stage. The model recovery ($\S$3.2) and self-induced labelling ($\S$3.3) dominate the offline stage, taking 40.4/26.3 hours. The reason the latter requires tremendous time is due to the lower GPU utilization caused by interactions with the score-head and sub-models.
@@ 5.5. Sensitivity analysis
[…]
Configuration sensitivity. We have the following conclusions in Figure 18. (1) 40M parameters are already a sweet configuration of TLM scale. A larger TLM’s gain on scoring and decision-making accuracy quickly gets eliminated by the overhead. (2) 12 shared bottom layers are reasonable for TLM, since more will lead to a smaller capacity of the heads, and less will incur a higher inference/training overhead. (3) A step size of 10% for model and prompt elastification is fine-grained enough for serving the diversified SLOs. Shrinking it to 5% makes almost no difference.
Ablation study. We show the effectiveness of ElastiLM’s key designs in Figure 18. We report the accuracy under a given resource SLO. Specifically, we respectively remove the key designs from our full system. Each removal of key design shows a clear performance degradation. The most significant loss comes from removing the reordering technique, showing over 10% accuracy loss. The reason is mainly due to that the reordering preserves the most important permutation consistent units. The score-head shows the least degradation, yet it is still obvious (about 5%). In a nutshell, the design of neuron-reordering, LoRA recovery and each head of the TLM are all non-trivial.
Varying real-world SLOs. In Figure 18 we report the delta accuracy of ElastiLM compared to its strongest baseline PFS (without considering switching overhead) on ARC_E LLaMA-7B. ElastiLM is close to or much better than PFS on all SLOs, especially those are more common on mobile LLM tasks (e.g., chatbot, UI automation agent or screen recorder).
@@
Quantization. Quantization is a widely-used technique for deploying LLMs on mobile devices. Here we employ a linear and weights only method on the $\alpha=0$ trace in Figure 19. ElastiLM can deliver LLM service with almost lossless accuracy under 8bits integers, and acceptable (3% lower absolute accuracy) under 4bits.
Concise prompts. We further emulate the scenario where the prompts have been consciously streamlined by the LLMaaS callers. We use LLMLingua2 to filter out about 15% verbose tokens in the prompt. The results (on trace with $\alpha$=0) in Figure 19 show that ElastiLM can still significantly outperform its baselines.
5.6. Discussion
@@
Generalization to other SLOs. Although in this paper we identify LLM inference latency as the SLO of requests, the SLO can also been easily generalized to other metrics with our proposed method. Here we define another two SLOs: end-to-end request inference time $SLO_{time}$ (i.e., prefill + decode stage) and inference energy consumption $SLO_{energy}$, each of which is with a step size of 10%. The results are obtained on Mi 14 smartphone. As shown in Figure 20, ElastiLM consistently outperforms its baselines on these new SLOs. the rationale is that these SLOs are all a specific kind of resource requirement and can be break down into prompt- and model- level elastifications. We believe that ElastiLM can serve various metrics of SLOs for diverse apps’ requirements.
Apps with competing demands. In regular on-device LLM service model, there is no concurrent requests since the device only has one user. However, in some extreme scenarios, other apps might have competing demands. ElastiLM will make it transparent to developers/apps. For instance, the LLM service may introduce upper-layer scheduling mechanisms like batching (Yu et al., 2022) to handle concurrency.
@@
Figure 21. “What if” experiment on mobile accelerators that cannot seamlessly run ElastiLM and its baselines currently. LG2+CS is omitted due to its sparse compute.
On mobile SoC accelerators. To the best of our knowledge currently (Aug., 2024) none of mobile DSAs can seamlessly run prefill and decode of ElastiLM and all its baselines due to legacy SDK/hardware issues like dynamic shape, graph building or non-sparse kernels. Thus, we conduct a “what if” experiment by mapping the theoretical compute/memory load to the profiled latency on these processors. We use MNN (Jiang et al., 2020) and mllm-NPU (Xu et al., 2024) for GPU/NPU profiling on Mi 14, respectively. The results are shown in Figure 21. ElastiLM can still significantly outperform the baselines. We believe that ElastiLM will become more friendly and practical to practitioners with the maturity of accelerators.
@@ 6. Related Work
Elastic neuron networks. Elastic neuron networks can change their capacity at runtime to dynamically adapt to various resource/accuracy constraints. Early exit networks (Teerapittayanon et al., 2017; Tambe et al., 2021) only perform inference on bottom layers since they are empirically more critical. Yet, early exit is not suitable for elastic LLM service. On one hand, due to LLMs’ autoregressive inference nature, a skipped layer’s KV cache may be accessed later. On the other hand, using layers as the granularity for trade-offs is not fine-grained enough. Parameter sharing networks (Fang et al., 2018; Han et al., 2021; Wen et al., 2023) generate memory-efficient sub-models that shares parameters with each other. For instance, NestDNN (Fang et al., 2018) and LegoDNN (Han et al., 2021) create sub-models of CNNs via offline re-pretraining, which is costly for foundation models. Adaptivenet (Wen et al., 2023) employs a CNN-oriented supernet, which provides diverse accuracy-latency trade-offs yet needs extensive pre-training and memory resources. Activation sparsity (Liu et al., 2023; Song et al., 2023; Cao et al., 2019; Kong et al., 2023) elasticitizes DNNs via sparsifying weights according to the NN inputs. An accuracy-latency trade-off can be achieved by setting the proportion of activated weights. However, the prefill stage cannot be elastictized due to the low locality.
@@
Efficient on-device LLM inference. Tremendous work (team, 2023; Jiang et al., 2020; mll, 2024; Xue et al., 2024) shed light on resource-efficiently deploying LLMs on mobile devices. For instance, MLC-LLM (team, 2023) is an NN-compiler with operator- and kernel- level optimizations for mobile devices. MNN (Jiang et al., 2020) and mllm (mll, 2024) are on-device inference libraries for LLMs. PowerinferV2 (Xue et al., 2024) addresses the memory issue of mobile devices by introducing swapping and activation sparsity to LLM inference. Targeting at elastic LLM service, ElastiLM is orthogonal to these work.
@@
Foundation models as a service. As a general task solver, or so-called “AGI”, a single foundation model is deployed as a service to process heterogeneous tasks (Yin et al., 2024; Yuan et al., 2024; Ge et al., 2023). For instance, AIOS (Ge et al., 2023) is a system where a single LLM runs as the “brain” to serve all apps on the device. ElastiLM makes it more practical on resource-constrained mobile devices.
@@
Model collaboration. ElastiLM employs a dual-head tiny language model to elasticize the prompt of the LLM. Using a small model to collaborate with the big model is common in ML systems. For instance, speculative decoding (Miao et al., 2024; Xu et al., 2023) accelerates the decode stage with a draft SLM. However, the prefill stage cannot be accelerated since it is typically compute-bound. LLMlingua (Jiang et al., 2023) uses an SLM to refine the LLM prompt, which has several untackled issues for elastic LLM service as discussed in this paper.
@@ 7. Conclusion
This work has proposed ElastiLM, an elastic on-device LLM service that serves apps with diverse SLOs. ElastiLM incorporates two novel designs, i.e., one-shot reordering of permutation consistent units and dual-head tiny language model to fully unleash the potential of model- and prompt- elastification. ElastiLM significantly outperforms competitive baselines by up to 14.83% and 10.45% on average in accuracy.
8. Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant 62325201. Mengwei Xu was supported in part by the Shenzhen Science and Technology Program with Grant No. JCYJ20241202124021028.